# When Federated Learning Meets Quantum Computing: Survey and Research Opportunities

### v2

@@ 9.1 Experimental Details
Upon analyzing the experimental evaluations in the literature, we summarize some crucial details in Table VI. MNIST [149] is observed as the most frequently used dataset, appearing in roughly one-third of studies; in contrast, specialized medical datasets (chest X-rays, MRI scans, etc.) represent an emerging application frontier. This preference for established benchmarks like MNIST suggests both practical convenience and a baseline foundation for comparability, even though it potentially limits exploration of quantum advantages in complex environments. FedAvg [1] dominates as the aggregation method, employed in approximately 70% of implementations, reflecting its simplicity and effectiveness. However, it possibly constrains exploration of quantum-specific aggregation techniques. The accuracy metrics span impressively from modest 52% (in non-IID mini-MNIST scenarios) to stellar 99% achievements, with medical applications consistently surpassing 90%. Yet these values resist meaningful comparison due to inconsistent experimental setups. Notably, only about 40% of studies address non-IID data characteristics, leaving a critical gap given that data heterogeneity represents a fundamental challenge in real-world FL scenarios. Additionally, in Fig. 12, we report the utilization of qubits and model accuracy by different datasets commonly used in QFL frameworks.
@@
Client participation structures vary dramatically from minimal setups with 2–3 clients to expansive networks with 100. Most studies cluster around modest client counts from 5 to 15, suggesting computational limitations when scaling quantum circuits in federated environments. Training rounds show a similar variation (spanning from 15 to 100 rounds), indicating a practical balance between computational feasibility and model convergence. This pattern reveals a field that is still in its infancy between theoretical exploration and practical implementation constraints. The studies show increased attention to adversarial robustness (with dedicated training regimens requiring fewer rounds) and specialized applications like power systems and speech emotion recognition. Some studies achieve an impressive accuracy of 98% with minimal client participation (3–5), while others could not exceed 52% despite extensive training.
From an implementation aspect, the utilization of various libraries highlights the quantum ecosystem’s maturation, with PennyLane [145] emerging as the dominant framework (appearing in at least 30% of specified implementations), followed by Tensorflow Quantum [147] and Qiskit [138]. This technological evolution from theoretical proposals to practical implementations facilitates reproducibility. The usage of D-Wave quantum annealer [182] in recent studies suggests growing interest in hardware-specific implementations beyond gate-based quantum computing. The striking accuracy disparity between conventional and quantum approaches on identical datasets raises important questions about quantum advantage in FL contexts. Are these differences attributable to quantum properties, or do they merely reflect varying classical pre-processing techniques? This ambiguity runs through the field, highlighting the need for more rigorous ablation studies and standardized evaluation frameworks to separate genuine quantum contributions from implementation artifacts. Despite promising results, the lack of standardized benchmarking protocols remains a significant obstacle to meaningful cross-study comparisons and reliable assessment of quantum advantages in FL.
@@
Finally, in Fig. 13, we illustrate the Gate fidelity [191] along with qubit accuracy with varying number of qubit states. The gate fidelity quantifies how well the quantum gate performs what it is supposed to do, ideally without any errors. When the fidelity reaches 100%, it implies the quantum gates are performing their intended operations without any errors, thus behaving ideally. For comparison, each box consists of the number of qubits utilized by the system. This comprehensive metric is vital since, despite high accuracy in individual qubits, low gate fidelity can accumulate errors throughout computations, thereby compromising the overall efficacy of quantum algorithms.
@@ 9.2 Major Findings
  • Several studies highlight that non-IID data significantly hinders achieving high accuracy and stable convergence. For instance, [64] shows a mere 81% accuracy after 20 rounds in non-IID settings, underscoring the difficulty of integrating varied client updates efficiently.
  • The introduction of quantum elements into classical FL strategies has proven advantageous. Studies such as [77] and [52] provide compelling evidence that incorporating quantum circuits with classical optimization tactics (like FedAvg) enhances performance. Remarkably, [52] achieves 99% accuracy on MNIST using only 6 clients over 200 rounds, indicating that fewer clients suffice when iterative refinement is employed.
  • Privacy preservation techniques are pivotal in these analyses. Quantum security approaches, as illustrated in [103][104], and [105], succeed in maintaining high accuracy with minimal computational costs. For example, [103] mentions a 20% reduction in overhead without compromising accuracy. These findings suggest that robust privacy implementations, when well-integrated, need not degrade efficiency.
  • Large-scale studies like those in [89] and [77] prove that QFL systems can scale from 10 to 100 clients while sustaining accuracy rates in the upper 90% range, showcasing the system’s ability to manage diverse client bases under challenging conditions. Emerging models, including the federated QLSTM in [61] and the quantum fuzzy neural network in [60], expand QFL’s reach beyond simple classification to include temporal data and fuzzy logic integration, potentially benefiting areas such as sensor networks and control systems. While these models may lack specific accuracy metrics, their swift convergence and adept handling of complex data suggest promising practical applications.
@@ 10 Discussion and Recommendations
QC has the potential to revolutionize FL by addressing its core challenges and enabling novel applications across diverse domains. The inherent parallelism of quantum systems facilitates exponential speedups in solving complex optimization problems, such as those encountered in model training and aggregation, which are computationally intensive in classical FL systems. This advantage is particularly impactful in fields like drug discovery and material science, where QFL could enable faster and more accurate predictions on quantum-specific data. Furthermore, the integration of quantum cryptographic techniques, such as QKD, ensures unprecedented data privacy and security, making FL more viable in sensitive sectors like healthcare and finance. The ability of quantum devices to process high-dimensional data and operate effectively in noise-prone environments opens pathways for applications in decentralized sensing networks, autonomous systems, and large-scale industrial IoT platforms. As quantum hardware matures and hybrid quantum-classical systems become more practical, QFL frameworks can leverage the advancements to achieve greater scalability, robustness, and efficiency, driving innovation in AI and beyond.
@@
$\bullet$ Quantum-enhanced dynamic server aggregation protocol: Centralized aggregation vulnerabilities in QFL require a new approach to server and client selection. Researchers can develop a quantum-probabilistic framework utilizing quantum randomness and quantum interference to ensure secure and unpredictable aggregation. By employing a quantum circuit for cryptographically secure random selection vectors, such a framework would support a multi-server protocol where nodes are chosen dynamically using quantum entropy, which can further address server collusion and client selection bias, thus forming a resilient distributed learning system by avoiding centralized aggregation weaknesses.
@@
$\bullet$ Quantum TL with quantum generative adversarial networks: Knowledge transfer in quantum computing is limited by architectural differences. Using quantum generative adversarial networks (QGANs) [192, 193], transferable feature representations can be created for different architectures. By training generator-discriminator pairs to translate feature distributions, a flexible TL framework can be developed. Such a system with quantum variational autoencoders and GANs can help establish a universal quantum feature translation mechanism, thereby making it a potential venue for further research.
$\bullet$ Quantum-resistant backdoor attack mitigation framework: Backdoor attacks pose a serious threat to QFL systems by injecting malicious training data that compromise the integrity of the model, which demands a quantum-enabled framework for detecting and mitigating these attacks, using quantum state tomography and interference patterns to probabilistically identify and neutralize threats. Such a framework would employ a quantum circuit that can analyze multiple feature spaces to detect anomalous patterns, offering a novel way to identify distributed attack vectors. It can shift backdoor detection from static pattern recognition to a dynamic quantum-assisted anomaly detection system, thus highly required for a sustainable QFL system.
$\bullet$ Categorization of security against quantum attacks, noise, and data privacy mechanisms: QFL introduces vulnerabilities at the intersection of quantum attacks, noise propagation, and data privacy. The lack of a unified framework to categorize and address these issues leads to fragmented and inefficient security solutions, limiting scalability and robustness. A solution would be to develop a multi-layered categorization framework that can systematically address these core challenges. This framework can use quantum-secure protocols, such as lattice-based cryptography, to mitigate quantum threats, while incorporating advanced error correction techniques.
$\bullet$ Quantum contextual invariant robustness framework: Distributed learning systems face challenges in maintaining consistency across varied computational environments. Our critical analysis offers a future direction to design a quantum mechanism that would leverage symmetry principles to stabilize FL by preserving core insights. Such a quantum circuit probabilistically can preserve contextual transformations, revolutionizing computational transfer management.
$\bullet$ Quantum entanglement-enhanced secure FL: Distributed ML systems are prone to privacy breaches. Developing a secure aggregation protocol using QKD and entangled qubit states for superior cryptographic security, would be a practical solution. By encoding model updates in quantum states, this protocol can ensure privacy at a quantum level. With QKD networks integrated into FL, one can achieve communication with privacy beyond classical cryptography.

### v4

@@ 9.1 Experimental Details
Upon analyzing the experimental evaluations in the literature, we summarize some crucial details in Table 9. MNIST [lecun1998gradient] is observed as the most frequently used dataset, appearing in roughly one-third of studies; in contrast, specialized medical datasets (chest X-rays, MRI scans, etc.) represent an emerging application frontier. This preference for established benchmarks like MNIST suggests both practical convenience and a baseline foundation for comparability, even though it potentially limits exploration of quantum advantages in complex environments. FedAvg [mcmahan2017communication] dominates as the aggregation method, employed in approximately 70% of implementations, reflecting its simplicity and effectiveness. However, it possibly constrains exploration of quantum-specific aggregation techniques. The accuracy metrics span impressively from modest 52% (in non-IID mini-MNIST scenarios) to stellar 99% achievements, with medical applications consistently surpassing 90%. Yet these values resist meaningful comparison due to inconsistent experimental setups. Notably, only about 40% of studies address non-IID data characteristics, leaving a critical gap given that data heterogeneity represents a fundamental challenge in real-world FL scenarios. Additionally, in Fig. 13, we report the utilization of qubits and model accuracy by different datasets commonly used in QFL frameworks.
@@
Client participation structures vary dramatically from minimal setups with 2–3 clients to expansive networks with 100. Most studies cluster around modest client counts from 5 to 15, suggesting computational limitations when scaling quantum circuits in federated environments. Training rounds show a similar variation (spanning from 15 to 100 rounds), indicating a practical balance between computational feasibility and model convergence. This pattern reveals a field that is still in its infancy between theoretical exploration and practical implementation constraints. The studies show increased attention to adversarial robustness (with dedicated training regimens requiring fewer rounds) and specialized applications like power systems and speech emotion recognition. Some studies achieve an impressive accuracy of 98% with minimal client participation (3–5).
From an implementation aspect, the utilization of various libraries highlights the quantum ecosystem’s maturation, with PennyLane [pennylane] emerging as the dominant framework (appearing in at least 30% of specified implementations), followed by Tensorflow Quantum [tfq] and Qiskit [qiskit]. This technological evolution from theoretical proposals to practical implementations facilitates reproducibility. The usage of D-Wave quantum annealer [dwave] in recent studies suggests growing interest in hardware-specific implementations beyond gate-based quantum computing. The striking accuracy disparity between conventional and quantum approaches on identical datasets raises important questions about quantum advantage in FL contexts - are these differences attributable to quantum properties, or do they merely reflect varying classical pre-processing techniques? Despite promising results, the lack of standardized benchmarking protocols remains a significant obstacle to meaningful cross-study comparisons and reliable assessment of quantum advantages in FL.
@@
Existing experimental studies on QFL studies are limited in scope, making statistical reliability crucial. Early frameworks validate feasibility with few clients and hybrid models[QFL-TL-2]. Other works show that most QFL frameworks still operate with limited participants, shallow parametrized quantum circuits, synthetic partitions, and idealized conditions, leaving open questions about the necessary client count and conditions for statistical reliability in varied settings [QFL-type, ballester_quantum_2025]. FL literature further indicates that client count, participation rate, and non-IID structure affect convergence and robustness highlights issues like label skew impacting simple aggregation techniques (such as FedAvg) [FL1, hsu_non_iid_2019]. Recent reviews on data challenges in FL further emphasize that unreliable participation, skewed sample sizes, and unreported data distributions can cause overly optimistic results that fail to generalize and address heterogeneity [saeed_datachallenges_2025, guendouzi_systematic_2023]. QFL inherits these issues and adds new challenges like shot noise, gate noise, device-specific error profiles, which together widen confidence intervals on performance metrics if the federation operates with too few or too homogeneous clients [QFL-TL-2, ballester_quantum_2025]. Current QFL experiments typically instantiate with several clients, often with synthetic non-IID splits and simplified quantum models limited by NISQ hardware limits. This choice is understandable from an implementation standpoint but it implies that many reported results are more feasibility demonstrations than proof of robustness at production scale.
Few studies measure changes in accuracy, variance, and robustness to adversaries change as the number of participating nodes increases or as non-IID severity intensifies [QFL-type, ballester_quantum_2025]. This is more challenging in QFL because quantum noise can mask or mimic heterogeneity effects, and small sample sizes on each device make variance estimates less stable. A better approach would be to consider the minimum number of clients as a derived quantity that depends on heterogeneity and noise, not as an arbitrary hyperparameter. Reliability should focus on variance across clients, reporting local metric dispersion and how it changes with different sampling strategies. Classical FL shows increasing clients improves risk stability under non-IID settings [FL1, hsu_non_iid_2019]. For QFL, this suggests that studies claiming strong generalization under heterogeneous conditions should (i) evaluate on configurations with numerous clients, (ii) use established non-IID protocols, and (iii) conduct multiple runs across random seeds to reveal outcome variations.
Evaluations should report confidence bands, partitions, and device noise realizations to expose the spread of outcomes. Rather than focusing solely on point estimates ofnot just accuracy or loss, to highlight the effects of limited client counts and quantum noise. Non-IID data generally interacts with QFL-specific constraints in subtle ways. Strong label skew or feature bias amplifies the influence of a few high-capacity quantum clients if their updates dominate aggregation, particularly when these clients host deeper parametrized circuits or larger shot budgets. Hybrid QFL frameworks that adopt nested ansatz families or masked aggregation strategies mitigate this imbalance by sharing a common parameter core across all clients and restricting deeper blocks to capable devices, yet this design choice needs to be validated empirically with metrics that capture fairness and influence distribution across clients [QFL-TL-2, FL1, ballester_quantum_2025]. Without such analysis, there is a risk that performance gains reported in small or symmetric testbeds conceal brittle behaviour under realistic deployment conditions where client capabilities and data distributions differ significantly.
Finally, along with ensuring practical environments for experiments, it is also to measure and analyze how well qubits have been utilized in a system, and what is their performance overall. In Fig. 14, we illustrate the Gate fidelity [borgarino2025demystifying] along with qubit accuracy with varying number of qubit states. The gate fidelity quantifies how well the quantum gate performs what it is supposed to do, ideally without any errors. When the fidelity reaches 100%, it implies the quantum gates are performing their intended operations without any errors, thus behaving ideally. For comparison, each box consists of the number of qubits utilized by the system. This comprehensive metric is vital since, despite high accuracy in individual qubits, low gate fidelity can accumulate errors throughout computations, thereby compromising the overall efficacy of quantum algorithms.
@@ 9.2 Major Findings
  • Several studies highlight that non-IID data significantly hinders achieving high accuracy and stable convergence. For instance, [QFL-QAOA] shows a mere 81% accuracy after 20 rounds in non-IID settings, underscoring the difficulty of integrating varied client updates efficiently.
  • The introduction of quantum elements into classical FL strategies has proven advantageous. Studies such as [QFL-TL-2] and [QFL-QNGD] provide compelling evidence that incorporating quantum circuits with classical optimization tactics (like FedAvg) enhances performance. Remarkably, [QFL-QNGD] achieves 99% accuracy on MNIST using only 6 clients over 200 rounds, indicating that fewer clients suffice when iterative refinement is employed.
  • Privacy preservation techniques are pivotal in these analyses. Quantum security approaches, as illustrated in [QFL-DPFHE][QFL-DPNoise], and [QFL-BlindC], succeed in maintaining high accuracy with minimal computational costs. For example, [QFL-DPFHE] mentions a 20% reduction in overhead without compromising accuracy. These findings suggest that robust privacy implementations, when well-integrated, need not degrade efficiency.
  • Large-scale studies like those in [QFL-QCNN] and [QFL-TL-2] prove that QFL systems can scale from 10 to 100 clients while sustaining accuracy rates in the upper 90% range, showcasing the system’s ability to manage diverse client bases under challenging conditions. Emerging models, including the federated QLSTM in [QFL-QLSTM] and the quantum fuzzy neural network in [QFL-QFuzzyNN], expand QFL’s reach beyond simple classification to include temporal data and fuzzy logic integration, potentially benefiting areas such as sensor networks and control systems. While these models may lack specific accuracy metrics, their swift convergence and adept handling of complex data suggest promising practical applications.
@@ 10 Discussion and Recommendations
QC has the potential to revolutionize FL by addressing its core challenges and enabling novel applications across diverse domains. The inherent parallelism of quantum systems facilitates exponential speedups in solving complex optimization problems, such as those encountered in model training and aggregation, which are computationally intensive in classical FL systems. This advantage is particularly impactful in fields like drug discovery and material science, where QFL could enable faster and more accurate predictions on quantum-specific data. The ability of quantum devices to process high-dimensional data and operate effectively in noise-prone environments opens pathways for applications in decentralized sensing networks, autonomous systems, and large-scale industrial IoT platforms. As quantum hardware matures and hybrid quantum-classical systems become more practical, QFL frameworks can leverage the advancements to achieve greater scalability, robustness, and efficiency, driving innovation in AI and beyond.
While the addition of QC to QFL brings numerous advantages, it also generates qualitatively new concerns that researchers must consider while building novel QFL frameworks. QFL emphasizes enhanced model inversion via quantum expressivity, allowing parametrized quantum circuits to offer richer hypotheses than classical models in certain tasks. This expressivity, combined with structured measurement statistics, can enhance the link between updates and client data. An adversary equipped with quantum or high-performance classical solvers may run reconstruction attacks, motivating attack risks from quantum-capable adversaries, warranting robust security definitions [zhao_federation_2025, carletti_sok_2025]. QFL also introduces threats like quantum channel and entanglement abuse, where compromised coordinators might exploit quantum networks to manipulate client systems beyond intended protocols. Existing QFL taxonomies do not capture this nonlocal manipulation capability. Recent QFL surveys already emphasize that security analyses rarely model malicious entanglement or untrusted quantum switches, indicating an open gap between protocol design and adversarial modeling [ballester_quantum_2025, nguyen_qfl_survey_2025].
Further, delegated QC in QFL merges cloud threats with FL assumptions, risking data integrity and confidentiality. A malicious quantum service can bias gradients, embed structured errors that evade simple norm-based anomaly detection, or harvest side-channel information about client circuits and data encodings. This behaviour blurs the boundary between model poisoning and hardware-level attacks and require joint analysis of delegated computation security and FL robust aggregation. The drive towards post-quantum cryptography and quantum-safe secure aggregation itself becomes a potential attack surface. Recent works on post quantum-secure FL aggregation demonstrate that careful parameterization is needed to avoid new leakage patterns or degradation in robustness [QFL-SecAggPostQ2, Zhang2025_SAReview]. In QFL, where all adversaries are, by construction, quantum-capable, flawed or partial deployments of post-quantum primitives can create a false sense of protection while leaving update streams exploitable.
On the threats aspect, while a broad spectrum of threats is conceivable in FL, early deployments of QFL necessitate a careful prioritization of those most likely to impair real-world pilots. The distinctive constraints of the NISQ era imply that certain attack vectors pose disproportionately higher risks [baseri2025futureproofingcloudsecurityquantum]. First, poisoning attacks on model updates and data inputs are a major concern. As in FL, even a few malicious updates can degrade the global model. This risk is greater in QFL due to limited qubits and high measurement variance, complicating the differentiation between benign noise and adversarial perturbations. Second, protecting client information through gradient updates or quantum states is crucial. Hybrid quantum-classical processing might expose sensitive data through transmitted gradients or reduced states. Third, protecting against untrusted computation nodes like cloud QPU providers or aggregation servers is critical. Honest-but-curious participants may try to deduce client contributions, emphasizing the need for quantum-secure multi-party protocols with minimal overhead. Other classes of threats, such as collusion between semi-honest nodes or gate parameter manipulation in large-scale quantum devices, are of medium to lower priority in the near term. While important for the long-term roadmap, their impact is constrained by current hardware capabilities and deployment scales. Addressing all of these with minimal overhead under NISQ constraints can form the foundation for secure and trustworthy QFL systems.
@@
$\bullet$ Dynamic server aggregation: Centralized aggregation exposes QFL to trust concentration and a single point of failure. A future direction could be a rotating server committee, selected in each round by randomness that is both unpredictable and publicly verifiable [HerreroCollantes2017QRNG, Liu2018DIQRNG]. Quantum random number generators and their device-independent variants can provide certified entropy, while key derivation functions transform that entropy into selection seeds. Adding these seeds generate auditable vectors that specify which servers and clients participate in aggregation, and their publication through cryptographic techniques allows independent verification of the process. Further, threshold secure aggregation must then operate across the rotating committee, with explicit parameters for quorum, dropout recovery, and rekeying during rotation [Bonawitz2017SecAgg, Bell2020SecAggPlus]. Designing such a framework distributes authority, curbs collusion, and introduces accountability, while also demanding systematic evaluation of its latency, privacy retention, and fairness relative to fixed-server baselines.
@@
$\bullet$ Quantum transfer learning with quantum generative adversarial networks: Knowledge transfer across quantum models is limited by architectural differences in encoders, ansätze, and readout, which leads to incompatible feature representations across devices. A practical way forward is to use quantum generative adversarial networks (QGANs) [dallaire2018quantum, ngo2023survey] to learn transferable features that bridge these mismatches. Concretely, one can train generator–discriminator pairs to translate measured feature distributions from a source architecture to a target one, so downstream tasks on the target benefit from the structure learned on the source. This translator can be coupled with a lightweight quantum variational autoencoder module to regularize the latent space and stabilize training, while the GAN objective focuses on distribution alignment. The resulting framework yields a flexible transfer pipeline: learn source features, map them through the QGAN translator to the target representation, then fine-tune on the target with minimal new data. Such a setup can support cross-device reuse in QFL where client hardware is heterogeneous, and opens a clear research track on stable, device-agnostic feature translation with rigorous evaluation under NISQ constraints.
$\bullet$ Quantum contextual invariant robustness framework: A persistent challenge in distributed learning is the degradation of stability when models operate across heterogeneous environments. Future research can address this by developing quantum circuits that exploit symmetry-preserving operations to maintain invariants during training and communication. Contextual transformations, when encoded as unitary symmetries, can be preserved through quantum contextuality, providing robustness against distortions induced by noise or cross-platform variability [Howard2014Contextuality, Raussendorf2013Contextuality]. Integrating these invariants with federated aggregation may allow models to retain consistency without retraining overhead, offering a structured pathway to improve cross-device transfer in QFL. The open question lies in quantifying how such invariants can be efficiently implemented on NISQ devices, where gate noise and limited qubits impose strict constraints on symmetry exploitation.
$\bullet$ Quantum entanglement-enhanced secure FL: Privacy in FL under classical secure aggregation is vulnerable due to collusion or side-channel leaks. A stronger method is to use the entanglement-assisted QKD, providing information-theoretic secure keys for encryption with one-time pads in every round, reducing exposure even with compromised relays [Ekert1991E91, Lo2012MDIQKD]. Entanglement allows measurement-device–independent protocols, eliminating reliance on trusted detectors, addressing loopholes in standard QKD [Scarani2009QKDRev, Pirandola2020QKDReview]. This maintains the classical nature of model updates while enhancing cryptographic strength [Wootters1982NoCloning]. While limitations include finite key rates, distance constraints, and reliance on underdeveloped quantum repeaters [Wehner2018QI], future work should assess partial-participation protocols under finite-key conditions and compare against forward-secure post-quantum cryptography to determine the benefits of entanglement-based integration.
$\bullet$ Quantum-resistant backdoor attack mitigation framework: Backdoor attacks remain a core risk in FL and extend to QFL through model replacement and distributed trigger injection [bagdasaryan2020how, xie2020dba]. Classical defenses rely on static anomaly detection, which often fails under distributed and adaptive poisoning strategies. A more promising direction is to combine robust aggregation with quantum subroutines that expose subtle correlations hidden from classical inspection. Swap-test circuits can compare suspect updates against trusted references to reveal anomalous overlaps [paris2004qse], while classical-shadow tomography enables efficient screening of many observables from few measurements [huang2020shadows]. These tools shift detection from coarse statistical checks to fine-grained quantum probes, providing resilience even in low-data regimes. The challenge remains to integrate them without excessive overhead on NISQ devices, motivating hybrid pipelines that use lightweight classical filters and escalate only confirmed cases to quantum-level scrutiny.
Acknowledgment
The work is supported by BITS Pilani Dubai Campus (BPDC) for the project “FaR-FedIoT: Fair and Resource-adaptive Federated System for IoT” under NFSG scheme.
[Uncaptioned image] Aakar Mathur is a senior undergraduate in Computer Science at BITS Pilani, Dubai Campus, UAE. He has worked on hands-on machine learning models at Emirates Airlines and has gained practical data analysis experience at Apparel Group. He has been working as a web development intern since 2025 at Siemens Energy, Dubai. His research interests include quantum computing, federated learning, machine & deep learning, and behavioural finance.
[Uncaptioned image] Ashish Gupta is an Assistant Professor in Computer Science at BITS Pilani Dubai Campus, UAE. He has been working as a postdoctoral fellow in the Department of Computer Science at Missouri S&T, USA, from 2021 to 2023. He received a Ph.D. in Computer Science and Engineering from the Indian Institute of Technology (BHU), Varanasi, India. His research interests include sensor data analytics, federated learning, and applied machine learning.
[Uncaptioned image] Sajal K. Das (Fellow, IEEE) is a Curator’s Distinguished Professor of Computer Science and the Daniel St. Clair Endowed Chair with Missouri University of Science and Technology, USA. His research interests include wireless and sensor networks, mobile and pervasive computing, mobile crowdsensing, cyber-physical systems and IoT, cloud and edge computing, cyber security, and social networks. He has co-authored more than 700 research articles in high-quality journals and refereed conference proceedings, five U.S. patents, and four books. His h-index is 100+ with more than 45,000 citations. He has been the Editor-in-Chief of Pervasive and Mobile Computing Journal (Elsevier’s) and an Associate Editor of IEEE Transactions on Dependable and Secure Computing.